OpenStudy (anonymous):

Sol each equation for 0°≤x<360°. 1-(1/4)tan(x)=3/4 Note that x is theta. Someone help me please.

OpenStudy (anonymous):

@Satan 

OpenStudy (dinamix):

@flatouthero tan(x) = 1 and argtan(x) = 45

OpenStudy (dinamix):

\[1-(1/4)*\tan(x) = 3/4 \] \[-(1/4)\tan(x) = (3/4)-1\] \[(-1/4) \tan(x) = -1/4\] \[\tan(x) = 1\] \[\tan^{-1} (1) = \frac{ \pi }{ 4 }+k \pi\] @TheProfessor35 @flatouthero( @satellite73 check my answer i think i am right

OpenStudy (anonymous):

looks right to me

OpenStudy (dinamix):

@satellite73 ty my friend

OpenStudy (theprofessor35):

@satellite73 you never communicate, I'm surprised!!